OpenStudy (anonymous):

Trigonometry Help. Find the exact values of the six trigonometric functions of given the point (-4, 5) on the terminal side of theta in standard position.

OpenStudy (anonymous):



OpenStudy (anonymous):

you need the length of the hypotenuse, which you find via pythagoras as \[h=\sqrt{4^2+5^2}=\sqrt{16+25}=\sqrt{41}\]

OpenStudy (anonymous):

then take the appropriate ratios for example \[\cos(\theta)=-\frac{4}{\sqrt{41}}\] etc

OpenStudy (anonymous):